OpenStudy (anonymous):

Y+x2=4x Y+4x=16 Find x and y

OpenStudy (mathmate):

Do you mean: \( Y+x^2=4x \\ Y+4x=16 \\ \) Subtract one equation from the other to get \( Y-Y3+x^2-4x = 4x -16\) Can you take it from here?

OpenStudy (mathmate):

Y3 should read Y.

OpenStudy (anonymous):

So Y3 is just y

OpenStudy (mathmate):

There was a typo, the last line should have read: \( Y-Y+x^2-4x = 4x -16\) which simplifies to: \( x^2-8x+16=0\) which is a quadratic you could solve by factoring. After that, you can evaluate Y from one of the two original equations.

OpenStudy (anonymous):

Ohhh okay thank you! :)

OpenStudy (mathmate):

yw!